OpenStudy (anonymous):

What is the average rate of change between x = 7 and x = 12 in the function below? x ; x < 5 f(x) = x + 2 ; 5 < x < 10 3x ;x > 10

OpenStudy (anonymous):

ANSWER CHOICES A. 5 2/5 B. 5 C. 27 D. 2 7/15

OpenStudy (anonymous):

Seems that I was wrong, haha.

OpenStudy (anonymous):

@bmp thanks for the help anyway but do yu know how to do this by the answer choices given?

OpenStudy (anonymous):

A. 5 2/5

OpenStudy (anonymous):

Avg = Δ y/ Δ x

OpenStudy (anonymous):

Δy = f ( 12) - f ( 7 ) / ( 12 - 7) = ( 36 - 9 ) / 5 = 27/ 5